OpenStudy (anonymous):

Compute this Bernoulli Trial: At least 4 failures in 7 trials with p = 0.75______?

Directrix (directrix):

Is p the probability of failure?

OpenStudy (anonymous):

success

Directrix (directrix):

4 failures: prob of failure is .25 prob of success is .75 C(7, 4) .25 ^ 4 times .75 ^ 3 =

Directrix (directrix):

5 failures C(7,5) .25^5 * .75 ^2 =

Directrix (directrix):

6 failures: C (7,6) .25 ^6 * .75 ^1 =

Directrix (directrix):

7 failures C(7,7) .25^7 * .75 ^0 = .25^7 =

Directrix (directrix):

@CW ---> Time for you to multiply and add

OpenStudy (anonymous):

sorry man i have been back and forth t

Directrix (directrix):

We have to do some arithmetic.

Directrix (directrix):

C(7, 4) .25 ^ 4 times .75 ^ 3 + C(7,5) .25^5 * .75 ^2 + C (7,6) .25 ^6 * .75 ^1 + C(7,7) .25^7 * .75 ^0 =

OpenStudy (anonymous):

I have my instructor on email now

OpenStudy (anonymous):

Thats why I am back and forth

Directrix (directrix):

Did you get .070556 approx?

Directrix (directrix):

Deal with your instructor first. Then, we'll finish this.

OpenStudy (anonymous):

He helped me answer a couple were done I am trying to do some supposed to be easier stuff with random variables and density functions

OpenStudy (anonymous):

I will post a couple more of these

OpenStudy (anonymous):

I have two more that i am stuck on